Solve and find particular solution of the first order non-linear differential equation of Bernoulli’s type with given condition.

(dy/dx)+(xy/1-x2 )=xy1/2 ; y(0)=1

 


1
Expert's answer
2021-02-24T06:49:11-0500

Solve and find particular solution of the first order non-linear differential equation of Bernoulli’s type with given condition.

(dy/dx)+(xy/1-x2 )=xy1/2 ; y(0)=1

Solution:

Divide the left and right sides of the equation by "\sqrt y" :

"\displaystyle\frac{1}{\sqrt y}\frac{dy}{dx}+\frac{x}{1-x^2}\sqrt y=x"

"\sqrt{y}=z" , "z'=\displaystyle\frac{y'}{2 \sqrt y}" , "\displaystyle\frac{y'}{\sqrt y}=2z'" .

"\displaystyle2z'+\frac{x}{1-x^2}z=x"

"z=uv" , "z'=u'v+v'u"

"\displaystyle2(u'v+v'u)+\frac{x}{1-x^2}uv=x"

"\displaystyle2u'v+u(2v'+\frac{x}{1-x^2}v)=x"

Let's compose and solve the system:

"\begin{cases}
 2v'+\frac{x}{1-x^2}v=0 \\
 2u'v=x 
\end{cases}"

From the first equation:

"2v'=-\frac{x}{1-x^2}v"

"2\frac{dv}{v}=-\frac{xdx}{1-x^2}"

"2\frac{dv}{v}=-\frac12\frac{dx^2}{1-x^2}"

"4\ln{|v|}=\ln{|1-x^2|}"

"v=(1-x^2)^{\frac14}"

Substitute "v" into the second equation:

"2u'(1-x^2)^{\frac14}=x"

"u=\displaystyle\int\frac{xdx}{2(1-x^2)^{\frac14}}=\displaystyle\int\frac{dx^2}{4(1-x^2)^{\frac14}}=-\frac13(1-x^2)^{\frac34}+C"

"z=uv=(1-x^2)^{\frac14}(-\frac13(1-x^2)^{\frac34}+C)=-\frac13(1-x^2)+C(1-x^2)^{\frac14}"

"\sqrt y=-\frac13(1-x^2)+C(1-x^2)^{\frac14}"

"y(0)=1"

"\sqrt 1=-\frac13(1-0^2)+C(1-0^2)^{\frac14}"

"1=-\frac13+C"

"C=\frac43"

"\sqrt y=-\frac13(1-x^2)+\frac43(1-x^2)^{\frac14}"

Answer: "\sqrt y=-\frac13(1-x^2)+\frac43(1-x^2)^{\frac14}" .